OpenStudy (anonymous):

f(x)=2x-1 and g(x)=x+2

OpenStudy (anonymous):

What are you supposed to do?

OpenStudy (shamil98):

no question here..

OpenStudy (anonymous):

find g o f

OpenStudy (anonymous):

replace g(x) with g(f(x)): g(f(x)) = f(x) + 2 g(f(x)) = 2x - 1 + 2 g(f(x)) = 2x + 1

OpenStudy (anonymous):

thank you